suppose x is a normally distributed random variable with μ = 50 and ϭ = 3. find a value of the random variable, call it x0, such that a. p(x ≤ x0) = 0.8413 b. p(x > x0) = 0.25 c. p(x > x0) = 0.95 d. p(41 ≤ x < x0) = 0.8630 e. 10% of the values of x are less than x0. f. 1% of the values of x are greater than x0.
1
Expert's answer
2020-08-03T18:27:08-0400

μ=50

σ=3


The Z score (Z) is obtained as:


"Z= (x- \u03bc)\/\u03c3"


Hence


"x=Z* \u03c3+ \u03bc"


Solution a)


"p(X \u2264 X_0 )=0.8413"


This is a left tail probability:


"p value=1-p(X \u2264 X_0 )=1-0.8413=0.1587"


At "p value=0.1587"

"Z score=-1"


Therefore:

"x=(-1)* 3+ 50"


Answer: 47.00


Solution b)


"p(X> X_0 )=0.25"


This is a right tail probability:

"p value=1-p(X \u2264 X_0)=1-0.25=0.75"


At "p value=0.75"


"Z score=-0.67"


Therefore:

"x=(-0.67)* 3+ 50"


Answer: 47.99


Solution c)

"p(X> X_0 )=0.95"


This is a right tail probability:

"p value=1-p(X \u2264 X_0)=1-0.95=0.05"


At "p value=0.05"

"Z score=1.64"


Therefore:

"x=(1.64)* 3+ 50"

Answer:  54.92


Solution d)

"p(41 \u2264X \u2264 X_0 )=0.8630"


This is a two tail probability:

"p value=1-p(X \u2264 X_0)=1-0.8630=0.137"


Obtain the left-side rejection region "p(41 <X)"

"Z= (x-\u03bc)\/\u03c3= (41- 50)\/3 = -3"

 

At "Z value=-3"

"p(41 <X)= 0.00135"


Hence, the probability can be expressed as a right tail probability as follows:

"p(X\u2264 X_0\n\n)=(0.8630+0.00135)=0.86435"


"p value=1-p(X \u2264 X_0) =1-0.86435=0.13565"

 

At "p value=0.13565"

"Z score=1.10"


Therefore:

"x=(1.10)* 3+ 50"


Answer:  54.10


Solution e)

If 10% of values of X are less than X_0, then

"p(X> X_0 )=0.90"


This is a left tail probability:

"p value=1-p(X \u2264 X_0)=1-0.90=0.10"


At "p value=0.10"

"Z score=-1.28"


Therefore:

"x=(-1.28)* 3+ 50"

 

Answer = 46.16


 

Solution f)

If 1% of values of X are greater than X_0, then

"p(X< X_0 )=0.99"


This is a right tail probability:

"p value=1-p(X \u2264 X_0)=1-0.99=0.01"


At "p value=0.01"

"Z score=2.33"


Therefore:

"x=(2.33)* 3+ 50"

 

Answer:  56.99